# Friendship-paradox paradox: Do most people’s friends really have more friends than they do?

### v2

@@ I Introduction
In practice, the colloquial phrasing of the friendship paradox is often taken to express a majority claim such as “for most people, their friends have more friends than they do.” This interpretation presumes that a typical individual is locally less connected than the people around them. The classical paradox, however, makes no such statement. It concerns only comparisons of network-wide averages and offers no guarantee about what happens for most nodes, or even for any specific node. As the examples in this paper illustrate, it is entirely possible for the paradox to hold in its traditional, mean-based form while a majority of nodes are in fact more connected than their neighbors on average. Conversely, one may also find networks where many nodes appear locally disadvantaged despite the population averages suggesting otherwise. These discrepancies arise because local comparisons depend sensitively on the distribution of neighbor degrees around each node, not merely on global averages. This observation motivates the need to distinguish clearly between population-level statements and majority-type, node-level assessments, and it sets the stage for the systematic framework developed in the remainder of the paper.
@@ Acknowledgements.
This research was supported by the National Research Foundation of Korea (NRF) under the grant RS-2021-NR061247. This research was supported by the National Research Foundation of Korea (NRF) under the grant RS-2021-NR061247. The author appreciates Krzysztof Sienicki for pointing out a subtle but important technical issue regarding the relationship between the condition $h_{i}<1/2$ and the neighborhood median, which helped improve the clarity and accuracy of the presentation.

### v3

@@ I Introduction
In practice, the colloquial phrasing of the friendship paradox is often taken to express a majority claim such as “for most people, their friends have more friends than they do.” This interpretation implicitly assumes that a typical individual is locally less connected than the people around them. The classical friendship paradox, however, makes no such statement: it concerns only network-wide averages and offers no guarantee about what happens for most nodes, or even for any specific node. A few earlier works—most notably those introducing and analyzing the strong friendship paradox [6, 7, 8]—have pointed out that local majority-based comparisons can behave very differently from the traditional mean-based formulation. Yet the conceptual distinction between population-level averages and node-level majority relations remains underemphasized in much of the broader literature. As the examples in this paper illustrate, it is entirely possible for both classical FP inequalities to hold while a majority of nodes are in fact more connected than the mean of their neighbors, or conversely, for many nodes to appear locally disadvantaged even when global averages suggest otherwise. These discrepancies arise because local comparisons depend sensitively on the distributional shape of neighbor degrees around each node, not merely on global means. Recognizing this gap motivates the need for a clearer framework separating mean-based and majority-based notions of “having more friends,” which we develop in the remainder of the paper.
@@ V Discussion and Outlook
Notes.—After the first version of this paper was posted, the author was kindly informed that the median-based majority condition studied here has been previously introduced under the name “strong friendship paradox” in several earlier works [6, 7, 8]. These studies also emphasized the role of distributional shape in separating mean-based and median-based comparisons, and documented empirical settings in which median-based dominance is widespread. The author gratefully acknowledges these important precedents and encourages readers to view the present contribution as complementary: the goal here is to provide a unified mathematical formulation of the mean-based and median-based quantities, clarify their logical independence, and highlight conditions under which majority-type inequalities diverge from the classical FP.
@@ Acknowledgements.
This research was supported by the National Research Foundation of Korea (NRF) under the grant RS-2021-NR061247. This research was supported by the National Research Foundation of Korea (NRF) under the grant RS-2021-NR061247. The author appreciates Krzysztof Sienicki for pointing out a subtle but important technical issue regarding the relationship between the condition $h_{i}<1/2$ and the neighborhood median, which helped improve the clarity and accuracy of the presentation. The author also thanks Kristina Lerman for kindly pointing out earlier works on the strong friendship paradox and for helpful correspondence that helped situate the present results within the broader literature.